OpenStudy (anonymous):

Let f(x) = 2x+5 and g(x) = x^2-3x+2. Perform each function operation and then find the domain. 4f(x) + 2g(x)

OpenStudy (anonymous):

4f(x) = 4 (2x+5 ) = 8x + 40 2g(x) = 2 (x^2-3x+2) = 2x² - 6x + 4 => 4f(x) + 2g(x) = 2x² + 2x + 44 Which is quadratic equation always have domain R.

OpenStudy (anonymous):

@Chlorophyll Thank you. Can you check if my work is correct here too http://openstudy.com/study#/updates/4f4ddc01e4b019d0ebae038a

OpenStudy (anonymous):

Did you make sure your post completed?

OpenStudy (anonymous):

What do you mean

OpenStudy (anonymous):

The one by your link is incompleted, however the one you just posted is completed, so it's fine!

OpenStudy (anonymous):

it's incomplete? whats missing?